OpenStudy (anonymous):

Bill is able to save $35/week after working parttime and paying his expenses. These two formulas show his weekly savings A. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 B. f(n) = 35n I. Which one of these formulas show the sequence written recursively and which shows it written explicitly? Justify your explanations. II. Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. Explain what your table means. III. Use any formula of your choice to find f(40). Explain why you chose that method and what your answer means. Show your calculations IV. Given the seq

OpenStudy (anonymous):

Please help I am beyond confused @ranga @abdela25

OpenStudy (anonymous):

@hartnn

OpenStudy (abdela25):

use this rule A. f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 make a table n | f(n) ----------- 1 | 35 2 | ? you need to find f(2) the rule says, for n>1 (and here n is 2, so we use it) f(n) = f(1)+ f(n-1) replace n with 2: f(2) = f(1) + f(2-1) or f(2) = f(1)+f(1) from the table, we see f(1) is 35, so you get f(2) = 35+35 f(2) = 70 now fill in n=3, and find f(3) can you do that ?

OpenStudy (anonymous):

That's still confusing but I can try.

OpenStudy (anonymous):

Thanks anyways

OpenStudy (abdela25):

try it it looks hard but it really isnt anymore help msg me